OpenStudy (anonymous):

Can you find the number if distinguishable permutations of the letters SWEET? is it 120?

OpenStudy (anonymous):

it's fact 5/fact 2 i guess.....

jimthompson5910 (jim_thompson5910):

You're on the right track

jimthompson5910 (jim_thompson5910):

But...the two E's are going to slightly alter the answer

jimthompson5910 (jim_thompson5910):

So you're doing to divide 120 by 2! = 2*1 = 2 to correct for double counting.

OpenStudy (anonymous):

so 60?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you! :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

but for the LUCKY word, it will be 120

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

okay! :)

jimthompson5910 (jim_thompson5910):

since each letter is different in LUCKY and there are 5 letters, it's simply 5! = 5*4*3*2*1 = 120

OpenStudy (anonymous):

okay got it :)

OpenStudy (anonymous):

thanks!!

jimthompson5910 (jim_thompson5910):

sure thing